What are next 4 terms of -1,2,12,40?
Answer by rothauserc(4718)

a(n) = (2n-1) * 2^n, where n = 0, 1, 2, 3, 4, ...
:
a(4) = (2*4-1) * 2^4 = 112
a(5) = (2*5-1) * 2^5 = 288
a(6) = (2*6-1) * 2^6 = 704
a(7) = (2*7-1) * 2^7 = 1664
:
******************************
-1,2,12,40,112,288,704,1664
******************************
: